When 2.35g of Magnesium (OH)2 is added to 250.0mL of water the temperature of water raises from 20.5°C to 36.0°C. What’s the molar enthalpy of the solution?


1
Expert's answer
2021-07-02T16:06:01-0400

Moles of Mg(OH)2 dissolved:


"n = \dfrac{Mass}{Molar\ mass} = \dfrac{2.35g}{58.3197\ g/mol} = 0.04\ mol"


Heat released:

"Q = mass\ of\ water\ *\ water\ specific\ heat\ capacity\ *\ temperature\ change = 250g\ *\ 4.18 \dfrac{J}{g\ *\ ^oC}\ *\ (36.0\ ^oC\ -\ 20.5\ ^oC) = 16197.5 J"


Molar enthalpy of the solution:

"\Delta H = -\dfrac{Heat\ released}{Moles\ of\ the\ solute} = -\dfrac{16197.5\ J}{0.04\ mol} = -404937.5\ J/mol \approx -404.9 \dfrac{kJ}{mol}"